OpenStudy (anonymous):

Suppose now that you wanted to determine the den-sity of a small yellow crystal to confirm that it is sulfur. From the literature, you know that sulfur has a density of 2.07 g/cm 3 . How would you prepare 20.0 mL of the liquid mixture having that density from pure sam-ples of CHCl3and CHBr3? (Note:1 mL = 1 cm 3

OpenStudy (aaronq):

what are the densities of the pure samples you're mixing?

OpenStudy (anonymous):

CHCL3 (d= 1.492 g/ml) and CHBr3 ( d=2.890)

OpenStudy (anonymous):

this was the previous question Consider the following: you mix 10.0 mL of CHCl3 (d= 1.492 g/mL) and 5.0 mL of CHBr3(d= 2.890g/mL), giving 15.0 mL of solution. What is the density of this mixture?

OpenStudy (aaronq):

so you wanna mix these in such a way to get the overall density to 2.07 g/mL let x be the mL of CHCl3 and y be the mL of CHBr3, then we have: y+x = 20 mL 1.492x+2.89y=2.07* 20.0 mL solve these equations

OpenStudy (anonymous):

CHCL3= 11.73 mL and CHBr3= 8.268 mL

OpenStudy (aaronq):

yep!

OpenStudy (anonymous):

thank you! @aaronq

OpenStudy (aaronq):

no problem!